OpenStudy (amy0799):

if f(x) = 4/(x-2), find f'(x)

OpenStudy (anonymous):

i am guessing that this is the very beginning of calc and you have to do with by hand, not using any shortcuts am i right?

OpenStudy (amy0799):

yup

OpenStudy (anonymous):

too bad

OpenStudy (anonymous):

next week you will say \[f'(x)=-\frac{4}{(x-2)^2}\] in your head, but i guess we can take the steps (long and arduous steps)

OpenStudy (anonymous):

we will have to go nice an slow first write the definition, then use some algebra ok a lot of algebra it is 99% algebra

OpenStudy (anonymous):

did you get to \[\huge \lim_{h\to 0}\frac{\frac{4}{(x+h-2)}-\frac{4}{x-2}}{h}\]

OpenStudy (amy0799):

yes.

OpenStudy (anonymous):

ok now in order to spare some agony lets forget about the h in the denominator, and forget about the limit we will deal with those last, ok?

OpenStudy (amy0799):

ok

OpenStudy (anonymous):

so the algebra we need is to (carefully) do this subraction \[\frac{4}{x+h-2}-\frac{4}{x-2}\]

OpenStudy (amy0799):

(4(x-2)-4(x+h-2))/((x-2)(x+h-2) is that right?

OpenStudy (anonymous):

can you do this? leave the denominator in factored form i.e don't multiply out

OpenStudy (anonymous):

yes that is correct now carefully multiply out in the numerator and combine like terms

OpenStudy (amy0799):

(4x-4x-8+8-4h)/((x-2)(x+h-2))

OpenStudy (anonymous):

ok that is multiplying out what is left in the numerator?

OpenStudy (amy0799):

-4h

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

now recall there is an \(h\) in the denominator cancel it

OpenStudy (anonymous):

what is left?

OpenStudy (amy0799):

-4/((x-2)(x+h-2))

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

now 'take the limit as h goes to zero" which is a fancy way of saying erase that h what is left?

OpenStudy (amy0799):

-4/(x-2)^2

OpenStudy (anonymous):

as promised (see above)

OpenStudy (amy0799):

thank u so much! :D

OpenStudy (anonymous):

YW!